The intensity of light X, measured in foot-candles, varies inversely as the square of the distance from the light d. If X=4/5 foot-candles when the distance is 20 feet. Find the constant of varieation and then find the intensity when the distance is 16 feet.
K=___________________
Answer_______________

Answer by venugopalramana(3286)

The intensity of light X, measured in foot-candles, varies inversely as the square of the distance from the light d. If X=4/5 foot-candles when the distance is 20 feet. Find the constant of varieation and then find the intensity when the distance is 16 feet.
K=___________________
The intensity of light X, measured in foot-candles, varies inversely as the square of the distance from the light d. HENCE...
INTENSITY = X =K/D^2...WHERE K IS A CONSTNT AND D IS THE DISTANCE FROM LIGHT SOURCE.
If X=4/5 foot-candles when the distance is 20 feet.
4/5=K/20*20...K=20*20*4/5=320
Find the constant of varieation
K=320
and then find the intensity when the distance is 16 feet.
X=320/16*16=5/4 FOOT CANDLES.